h1 h2

Consider a water pipe with a restricted section as illustrated beside. The consumption meter from the utility company indicates a constant water flow of 480 L/min. You also know that the pipe has a diameter of 40 cm at the inlet and outlet but only a diameter of 20 cm in the restricted region.

• B) What is the water column height above the restricted section of the pipe (that’s h2)?

C) What is the highest value (in kPa) of the absolute pressure in this water pipe?

D) Assume that the highest flow rate which can be reached when using this water pipe occurs when the fluid velocity in the restricted section (that’s v2) is 1 m/s at which point the pipe starts to vibrate and might, over time, fissure. What is the maximum water flow rate (in m3/min) associated with this extreme mode of operation?
Answer #1

Given,

Flow rate (Q)= 480L/min

Dia of inlet and outlet (D) = 40cm =0.4m

Dia of restricted section (d)= 20cm = 0.2m

we know

Density of water = 1000kg/m3 = 1 kg/l

So, 1m3 of water = 1000L

so, flow rate in m3/sec = (480 * 10-3)/ 60

Q=0.008 m3/sec

Now,

B)

velocity in point 1

v_{1}=rac{Q}{A_{1}}

0.008 m/sec π D2

0.008 mse π*0.42

0.064m/sec

velocity at point 2

v_{2}=rac{Q}{A_{2}}

Un_ 0.008 2 m/sec

0.008 m/sec π *0.22

V2 0.255m /sec

By Bernoulli's Theorem, we get

P_{1}+rac{1}{2} ho v_{1}^{2}+P.E_{1}=P_{2}+rac{1}{2} ho v_{2}^{2}+P.E_{2}

as potential energy changes with the change of height. Here, height does not change so change in potential energy is also 0.

Therefore P.E1=P.E2

Equation becomes

2

Rightarrow P_{1}-P_{2}=rac{1}{2} ho v_{2}^{2}-rac{1}{2} ho v_{1}^{2}

Rightarrow rac{P_{1}-P_{2}}{ ho }=rac{1}{2}left ( v_{2}^{2}- v_{1}^{2} ight )

dividing acceleration due to gravity (g) from both sides we get

Rightarrow rac{P_{1}-P_{2}}{ ho g }=rac{1}{2g}left ( v_{2}^{2}- v_{1}^{2} ight )

Since, P= ho gh

Rightarrow h_{1}-h_{2}=rac{1}{2g}left ( v_{2}^{2}- v_{1}^{2} ight )

Rightarrow h_{1}-h_{2}=rac{1}{2g}left ( 0.255^{2}- 0.064^{2} ight )

h1- h2 -0.03m

Note: we cannot find the height h2 unless the pressure at the given point is given or pressure at point 1 is given. So, we can calculate only the difference in pressure which is equal to 0.03m

C)

P_{absolute}=P+P_{atm}

in order to get absolute pressure, pressure in the pipe or height of water column must be given

D)

Given,

v2 = 1m/s

to get flow rate

we know,

Q= AV

Rightarrow Q= rac{pi *d^{2}}{4}* 1 m^{3}/sec

Rightarrow Q= rac{pi *0.2^{2}}{4}m^{3}/sec

Rightarrow Q= 0.0314m^{3}/sec

Rightarrow Q= 0.0314*60m^{3}/min

herefore Q=1.884m^{3}/min

If any doubt feel free to comment